# Fine-Tuning Stable Diffusion XL for Stylistic Icon Generation: A Comparison of Caption Size

### v1

@@ II Methods
In this section we detail the approaches and techniques used to fine-tune Stable Diffusion XL in order to generate icons pertaining to a specific preset style for commercial use. We explore various aspects including inference steps, and the integration of different sized captions to align generated images with our specific icon style. Additionally, we compare the outcomes using the same captions with DALL-E 3 to understand the CLIP and FID score comparison which also shows how these metrics fluctuate relative to the actual quality of the icons assessed by a human.
Figure 1: Distribution of file types for icon images related to kitchen cabinets, kitchens, and screws in commercial data provided.
@@
For screw icons, we trained three versions of Stable Diffusion XL using the same set of images but with different prompts to see the variation in icon quality, FID score and CLIP score. The model variants were short prompts with class images, short prompts and long prompts. The short prompt with class images leveraged short keywords and 20 high-quality Home Depot head-type icon images. The short prompt contained short keyword prompts whereas the long prompt contained detailed descriptions of the training dataset images. A fourth model used for comparison purposes was the DALL-E 3, which was used to generate images using the same inference prompt. The purpose of this model was to understand performance metric variation relative to the quality by human assessment. The parameters used for each model can be seen in Table I.
Prior-preservation loss is effective in encouraging output diversity and in overcoming language-drift in ”DreamBooth” [9]. Therefore, we add class-specific prior preservation loss to the Stable Diffusion XL model with short prompt during training using class images. This approach requires additional class images and careful balancing between instance-specific and class-generic images. By default, the prior_loss_weight is set to 1. This parameter controls the weight of the prior preservation loss during training, which helps to preserve the generalization capabilities of the model by penalizing deviations from the class-specific distribution.
@@
For kitchen cabinet icons, we also trained three versions of Stable Diffusion XL using the same set of images but with different prompts to see the variation in icon quality, FID score and CLIP score. In terms of training data, since there are a limited number of kitchen cabinet icons provided by The Home Depot, we gather multiple free kitchen icons from the internet in a similar manner to our screws methodology which brings our final dataset to be 42 icons. Class images are a set of reference images used to guide the training process of a machine learning model. These images are used to help the model understand the style and characteristics of the icons we want to generate. In terms of class images, we leverage 100 icons of appliances and furniture from Home Depot, all with the style we’re hoping to generate.
@@ III-B Models trained on commercial data for Kitchen Cabinets
Similar to the findings from screw icons, the short prompt trained model has a higher CLIP score. However, when looking at the FID score, the short and long prompt have very similar FID scores as seen in Figure 8. DALL-E 3 has a better CLIP score 13 but this does not necessarily mean that the icon is what we are looking for, which makes sense. Although the short prompt trained model has better performance than the long prompt model, when looking at CLIP and FID scores, there are cases where long prompt generated icons are visually better and more customized based on the description in the prompt. This can be seen in Figure 10(b) and Figure 16(b). This further substantiates the fact that a high CLIP score is not a direct measure of whether a generated icon is in the correct style.
@@
Additionally, when comparing results from Stable Diffusion XL with DALL-E 3, we observe a noticeable difference in style. This highlights the importance of training data in generating customized and consistent styles. To explore this further, we trained several models with a combination of icons and realistic product photos as the training data and included style variable in the caption json file. The results show that incorporating even a few realistic product photos can significantly skew the generated icon style to match that of the realistic photos. This can be seen in Figures 11(c)11(b)11(a).
@@ V-C Models trained on non-commercial data for Screws
Figure 18: Example generated icons based on commercial training data using long prompts
@@
Figure 19: Example generated icons based on commercial training data using short prompts
@@ V-D An analysis of inference steps and file-name keyword only training
When testing the inference query on several different types of steps, it can be evident than an optimal point is required to find the best parameter value. As we increased the inference steps on a basic trained version of the model (sub-optimal) using keywords only for training, we found the image icons to get better as the inference steps increased. However there is a point where the inference steps being too high could cause a bad quality icon. The reasoning for the imperfections here are due to smaller trained captions, smaller than our short captions with only file-name parsed keywords 21.
When training the Stable Diffusion XL model on only keywords parsed from the file-names themselves, such as ”Phillips head” or ”hex screw”, the results are not as good, however are still somewhat coherent. This tells us that without a unique identifier such as our use of ”TOK”, it is still possible to get the shape correctly for some cases, however some cases provide non-coherent icons 20.

### v2

@@ II Methods
In this section we detail the approaches and techniques used to fine-tune Stable Diffusion XL in order to generate icons pertaining to a specific preset style for commercial use. We explore various aspects including inference steps, and the integration of different sized captions to align generated images with our specific icon style. Additionally, we compare the outcomes using the same captions with DALL-E 3 to understand the CLIP and FID score comparison which also shows how these metrics fluctuate relatively to the actual quality of the icons assessed by a human.
Figure 1: Distribution of file types for icon images related to kitchen cabinets, kitchens, and screws in the provided commercial data.
@@
For screw icons, we trained three versions of Stable Diffusion XL using the same set of images but with different prompts to see the variation in icon quality, FID score and CLIP score. The model variants were short prompts with class images, short prompts and long prompts. The short prompt with class images leveraged short keywords and 20 high-quality Home Depot head-type icon images. The short prompt contained short keyword prompts whereas the long prompt contained detailed descriptions of the training dataset images. A fourth model used for comparison purposes was the DALL-E 3, which was accessed via API to generate images using the same inference prompt. The purpose of including DALL-E 3 was to understand performance metric variation relative to the quality by human assessment. The parameters used for the three Stable Diffusion XL models can be seen in Table I.
Prior-preservation loss is effective in encouraging output diversity and in overcoming language-drift in ”DreamBooth” [9]. Therefore, we add class-specific prior preservation loss to the Stable Diffusion XL model with short prompt during training using class images. This approach requires additional class images and careful balancing between instance-specific and class-generic images. By default, the prior_loss_weight is set to 1. This parameter controls the weight of the prior preservation loss during training, which helps to preserve the generalization capabilities of the model by penalizing deviations from the class-specific distribution. Class images are a set of reference images used to guide the training process of a machine learning model. These images are used to help the model understand the style and characteristics of the icons we want to generate.
@@
For kitchen cabinet icons, we also trained three versions of Stable Diffusion XL using the same set of images but with different prompts to see the variation in icon quality, FID score and CLIP score. In terms of training data, since there are a limited number of kitchen cabinet icons provided by The Home Depot, we gather multiple free kitchen icons from the internet in a similar manner to our screws methodology which brings our final dataset to be 42 icons. In terms of class images, we leverage 100 icons of appliances and furniture from Home Depot, all with the style we’re hoping to generate.
@@ III-B Models trained on commercial data for Kitchen Cabinets
Similar to the findings from screw icons, the short prompt trained model has a higher CLIP score. However, when looking at the FID score, the short and long prompt have very similar FID scores as seen in Figure 8. DALL-E 3 has a better CLIP score (see Figure 13), but this does not necessarily mean that the icon is what we are looking for, which makes sense. Although the short prompt trained model has better performance than the long prompt model, when looking at CLIP and FID scores, there are cases where long prompt generated icons are visually better and more customized based on the description in the prompt. This can be seen in Figure 10(b) and Figure 16(b). This further substantiates the fact that a high CLIP score is not a direct measure of whether a generated icon is in the correct style.
@@
Additionally, when comparing results from Stable Diffusion XL with DALL-E 3, we observe a noticeable difference in style. This highlights the importance of training data in generating customized and consistent styles. To explore this further, we trained several models with a combination of icons and realistic product photos as the training data and included style variable in the caption jsonl file. The results show that incorporating even a few realistic product photos can significantly skew the generated icon style to match that of the realistic photos. This can be seen in Figures 11(c)11(b)11(a).
@@ V-C Models trained on non-commercial data for Screws
Figure 18: Example generated icons based on public training data using long prompts
@@
Figure 19: Example generated icons based on public training data using short prompts
@@ V-D An analysis of inference steps and file-name keyword only training
When testing the inference query on several different types of steps, it can be evident than an optimal point is required to find the best parameter value. As we increased the inference steps on a basic trained version of the model (sub-optimal) using keywords only for training, we found the image icons to get better as the inference steps increased. However there is a point where the inference steps being too high could cause a bad quality icon. The reasoning for the imperfections here ais due to smaller trained captions, smaller than our short captions with only file-name parsed keywords (see Figure 21).
When training the Stable Diffusion XL model on only keywords parsed from the file-names themselves, such as ”Phillips head” or ”hex screw”, the results are not as good, however are still somewhat coherent. This tells us that without a unique identifier such as our use of ”TOK”, it is still possible to get the shape correctly for some cases, however some cases provide non-coherent icons (see Figure  20).